# Generic consistency and nondegeneracy of vertically parametrized systems

### v3

@@ 2.5. Main implications on dimension, Zariski denseness and nondegeneracy
    For generic $p\in\mathbb{C}^{k}$, all zeros of $F_{p}$ in $(\mathbb{C}^{*})^{n}$ are nondegenerate.
@@
    For generic $p\in\mathbb{C}^{k}$, $\mathbb{V}_{\mathbb{C}^{*}}(F_{p})$ is a nonsingular complex algebraic variety.
@@
    (degAllG) implies (real), (rad) and (reg).
@@
For (ii), for the implication (setZ)$\Rightarrow$(dimG), Lemma 2.6 gives that $\mathcal{Z}_{\mathcal{P}}$ is Zariski dense whenever $\mathcal{Z}_{\mathcal{P}}(\mathcal{X})$ is, and hence (dimG) follows from Theorem 2.2(i). The implication (dimG)$\Rightarrow$(dim1) is immediate. Finally, if (dim1) holds, then Theorem 2.2 gives that $\mathcal{Z}_{\mathcal{P}}$ is Zariski dense, and by Lemma 2.6 so is $\mathcal{Z}_{\mathcal{P}}(\mathcal{X})$, giving (setZ).
@@
Finally, for (vi), the implications from (degAllG) to (real) and (reg) follow from Proposition 2.9. The implication (degAllG)$\Rightarrow$(rad) follows from Proposition 2.14. ∎
@@
is irreducible, Theorem 2.2 holds also in the affine setting over $\mathbb{C}$ (rather than $\mathbb{C}^{*}$) after replacing $\mathbb{V}_{\mathbb{C}^{*}}(F_{p})$ by $\mathbb{V}_{\mathbb{C}}(F_{p})$. Additionally, Lemma 2.6, Propositions 2.9, 2.12 and 2.14, as well as Theorem 2.13 extend easily to $\mathbb{C}$ by allowing $\mathcal{X}\subseteq\mathbb{C}^{n}$. We conclude that Theorem 2.15 holds in the affine setting over $\mathbb{C}$ as long as $\mathcal{I}_{\mathbb{C}}$ is irreducible and $\mathcal{I}_{\mathbb{C}}\cap(\mathcal{P}\times\mathcal{X})$ is Zariski dense in $\mathcal{I}_{\mathbb{C}}$.
In Section 3, we will see that for augmented vertically parametrized systems, the eight first statements (deg1)–(dimG) from the beginning of the subsection are equivalent. Thus, all the desired properties of the polynomial system will rely on the existence of a nondegenerate zero in the complex torus. The key part of the argument will be to show that $\mathcal{Z}$ has nonempty Euclidean interior if and only if $F$ has a nondegenerate zero. This phenomenon does not need to happen for general families, as example (1.8) shows.
@@ 2.6. Zariski denseness and nonempty Euclidean interior
For part (i), as $S$ is constructible, $S=\bigcup_{i=1}^{t}Z_{i}\cap U_{i}$, for some irreducible Zariski closed sets $Z_{i}\subseteq\mathbb{C}^{k}$ and nonempty Zariski open subsets $U_{i}\subseteq\mathbb{C}^{k}$. The statement now follows from the fact that the Zariski closure satisfies $\overline{S}=\bigcup_{i=1}^{t}Z_{i}$. In case (ii), [BCR98, Prop. 2.8.2] gives that $S$ is Zariski dense in $\mathbb{C}^{k}$ if and only if its semialgebraic dimension is $k$, which in turn is equivalent to $S$ having nonempty Euclidean interior in $\mathbb{R}^{k}$. ∎
@@ 3.3. Nondegenerate zeros and generic consistency
Let $G\in\mathbb{C}^{m\times(m-s)}$ be Gale dual to $C$, in the sense that the columns of $G$ are a basis for $\ker(C)$. The parametrization $\phi$ in (3.3) gives rise to a parametrization of $\mathcal{Z}$ via the Laurent polynomial map
@@
As $\mathcal{D}=\Phi(\Delta)$ by Proposition 3.2(iii), it follows from Proposition 3.3 that $\mathcal{D}$ is the set of critical values of $\Phi$. Now Sard’s Lemma (see, e.g., [SEDS17, Prop. B.2] and [Mum76, Prop. 3.7]), adapted to very affine varieties, gives that $\dim(\overline{\mathcal{D}})<m+\ell$. ∎
@@
For $\mathbb{G}=\mathbb{R}^{*}$ and $\mathbb{G}=\mathbb{R}_{>0}$, it is straightforward to verify that restricting the map $\phi\colon\ker(C)\times(\mathbb{C}^{*})^{n}\to\mathcal{I}$ to $(\ker(N)\cap\mathbb{G}^{m})\times\mathbb{G}^{n}$ provides a parametrization of $\mathcal{I}\cap((\mathbb{G}^{m}\times\mathbb{R}^{\ell})\times\mathbb{G}^{n})$, which after projection gives a parametrization of $\mathcal{Z}_{\mathbb{G}^{m}\times\mathbb{R}^{\ell}}(\mathbb{G}^{n})$.
@@ 3.4. The main theorem on generic dimension
  • The statements (deg1), (degX1), (degXG), (degAllG), (setE), (setZ), (dimG), (dim1) are all equivalent to the condition
@@
  • Any of the statements mentioned above implies (dimX), (real) and (reg).
@@
The equivalence between (deg1) and the matrix rank condition is Proposition 3.2(ii). As $\mathcal{I}$ is irreducible of dimension $m+n-s$ by Theorem 3.1, the assumptions for Theorem 2.15 are satisfied. Therefore, for the first two bullet points it suffices to show that (setZ) implies (degAllG) and (setE). The equivalence between (setZ) and (setE) is Proposition 2.20(ii). By Lemma 2.6 and Proposition 2.20, $\mathcal{Z}$ has nonempty Euclidean interior in $\mathbb{C}^{m}\times\mathbb{C}^{\ell}$ if (setZ) holds. By Theorem 3.5, this implies (degAllG). Finally, if (degAllG) holds, so does (rad) by Theorem 2.15. If, on the other hand, (degAllG) does not hold, then the first bullet point of the statement gives that (setZ) does not hold either, and hence it follows by Theorem 2.2 that $\mathbb{V}_{\mathbb{C}^{*}}(F_{p})=\varnothing$ for generic $(a,b)\in\mathbb{C}^{m}\times\mathbb{C}^{\ell}$. Thus, generically, $\langle F_{a,b}\rangle=\mathbb{C}[x^{\pm}]$, which is radical. ∎
@@
    If (degAllG) holds under the assumptions of Theorem 3.7, then we also have that all zeros of $F_{p}$ are nondegenerate for generic $p\in\mathcal{Z}_{\mathcal{P}}(\mathcal{X})$.
@@
    When $\ell=0$, the rank condition in the first bullet point of Theorem 3.7 simplifies to
@@
Theorem 3.7 tells us that for augmented vertical systems, only one of two extreme scenarios occurs, provided the parameter sets $\mathcal{A}$ and $\mathcal{B}$ and the domain $\mathcal{X}$ of the variables satisfy some mild conditions: Either the system is generically consistent, has generically the expected dimension $n-(s+\ell)$, and generically all zeros are nondegenerate, or the system is generically inconsistent, has never the expected dimension, and all zeros are degenerate.
@@
The conclusions of Theorem 3.7 also hold for parametric systems $F_{\cdot,b}$ given by the restriction of an augmented vertical system $F$ to a fixed value of $b\in L((\mathbb{C}^{*})^{n})$. In this case, the incidence variety is irreducible of dimension $n+m-(s+\ell)$, with parametrization
@@ 3.5. Computational considerations
The rank condition from Theorem 3.7 can be checked in the following way: Let $G\in\mathbb{C}^{m\times(m-s)}$ be a Gale dual to $C$, whose columns form a basis for $\ker(C)$. We then want to check whether there exists some $(u,h)\in\mathbb{C}^{m-s}\times(\mathbb{C}^{*})^{n}$ such that
@@ 3.6. Freely parametrized systems
As a corollary of our results, we recover a well-known consistency characterization for freely parametrized systems. In particular, we obtain an elementary proof of [Yu16, Lemma 1] and [Kho16, Theorem 11].
Given subsets $\mathcal{S}_{1},\dots,\mathcal{S}_{s}\subseteq\mathbb{Z}^{n}$, we let
@@
and say that a property holds generically in $\mathcal{F}_{\mathrm{free}}(\mathcal{S}_{1},\dots,\mathcal{S}_{s})$ if it does so under the isomorphism $\mathcal{F}_{\mathrm{free}}(\mathcal{S}_{1},\dots,\mathcal{S}_{s})\cong\prod_{i=1}^{s}\mathbb{C}^{\mathcal{S}_{i}}$ that identifies each polynomial with its coefficients. Let
$\operatorname{Lin}(\mathcal{S}_{i}):=\operatorname{span}_{\mathbb{C}}\{u-v:u,v\in\mathcal{S}_{i}\}\subseteq\mathbb{C}^{n}$
denote the direction of the affine hull of $\mathcal{S}_{i}$.
@@
Similarly, let $M=[M_{1}\ M_{2}\ \cdots\ M_{s}]\in\mathbb{Z}^{n\times m}$ be the block matrix where the columns of $M_{i}\in\mathbb{Z}^{n\times m_{i}}$ are the elements of $\mathcal{S}_{i}$ in some fixed order. For the vertical system $F=C(a\star x^{M})$, it holds that $\mathcal{F}_{\mathrm{free}}(\mathcal{S}_{1},\dots,\mathcal{S}_{s})=\{F_{a}:a\in\mathbb{C}^{m}\}$. We refer to this $F$ as the vertical system associated with $\mathcal{S}_{1},\dots,\mathcal{S}_{s}$.
@@
Let $\mathcal{S}_{1},\dots,\mathcal{S}_{s}\subseteq\mathbb{Z}^{n}$ and for $i=1,\dots,s$, let $\mathcal{G}_{i}\subseteq\mathbb{C}^{n}$ be a generating set of $\operatorname{Lin}(\mathcal{S}_{i})$. The following statements are equivalent:
@@
    There exists a linearly independent tuple of vectors $(v_{1},\ldots,v_{s})\in\prod_{i=1}^{s}\mathcal{G}_{i}$.
@@
    There exists a linearly independent tuple of vectors $(v_{1},\ldots,v_{s})\in\prod_{i=1}^{s}\operatorname{Lin}(\mathcal{S}_{i})$.
@@
$C_{i}\operatorname{diag}(G_{i}u_{i})=[u_{i1}\quad u_{i2}\quad\dots\quad u_{i,m_{i}-1}\quad-(u_{i1}+\dots+u_{i,m_{i}-1})].$
@@
The specialization of Theorem 3.7 for vertical systems to the case $\mathcal{A}=\mathbb{R}^{m}_{>0}$ and $\mathcal{X}=\mathbb{R}^{n}_{>0}$ allows us to study freely parametrized systems with real coefficients with prescribed sign. Specifically, we consider maximal dimensional (open) orthants $\mathcal{O}_{i}$ of $\mathbb{R}^{\mathcal{S}_{i}}$, and consider the subset $\mathcal{F}_{\mathcal{O}_{1},\dots,\mathcal{O}_{s}}(\mathcal{S}_{1},\dots,\mathcal{S}_{s})$ of $\mathcal{F}_{\mathrm{free}}(\mathcal{S}_{1},\dots,\mathcal{S}_{s})$ consisting of the polynomials in the image of $\prod_{i=1}^{s}\mathcal{O}_{i}$ under the isomorphism $\prod_{i=1}^{s}\mathbb{C}^{\mathcal{S}_{i}}\cong\mathcal{F}_{\mathrm{free}}(\mathcal{S}_{1},\dots,\mathcal{S}_{s})$.
@@
Let $\mathcal{S}_{1},\dots,\mathcal{S}_{s}\subseteq\mathbb{Z}^{n}$. For $i=1,\dots,s$, let $\mathcal{G}_{i}\subseteq\mathbb{C}^{n}$ be a generating set of $\operatorname{Lin}(\mathcal{S}_{i})$ and $\mathcal{O}_{i}$ be a maximal dimensional orthant of $\mathbb{R}^{\mathcal{S}_{i}}$. The following statements are equivalent:
@@
    The semialgebraic set $\mathbb{V}_{\mathbb{R}^{*}}(F)\cap\mathbb{R}^{n}_{>0}$ is nonempty and of dimension $n-s$ for $F$ in a subset of $\mathcal{F}_{\mathcal{O}_{1},\dots,\mathcal{O}_{s}}(\mathcal{S}_{1},\dots,\mathcal{S}_{s})$ with nonempty Euclidean interior.
@@
    There exists a linearly independent tuple of vectors $(v_{1},\ldots,v_{s})\in\prod_{i=1}^{s}\mathcal{G}_{i}$ and $\mathcal{O}_{i}\notin\{\mathbb{R}^{\mathcal{S}_{i}}_{>0},\mathbb{R}^{\mathcal{S}_{i}}_{<0}\}$ for all $i=1,\dots,s$.
@@
In the freely parametrized setting, it is known from [BS24, Prop. 1.12] that one can obtain the generic properties predicted by Corollary 3.14, by only requiring the coefficients corresponding to the vertices of the Newton polytopes to be generic (while the other coefficients can be arbitrarily fixed). It is an interesting problem for future research to determine which parameters can be fixed in a vertical system while still retaining the generic properties given by Theorem 3.7.

### v4

@@ 2.5. Main implications on dimension, Zariski denseness and nondegeneracy
    For generic $p\in{\mathcal{Z}}$, all zeros of $F_{p}$ in $(\mathbb{C}^{*})^{n}$ are nondegenerate.
@@
  • (flatG)
    The projection $\pi\colon\mathcal{I}\to\mathbb{C}^{k}$ is flat at generic $(p,x)\in\mathcal{I}$.
@@
    $\mathbb{V}_{\mathbb{C}^{*}}(F_{p})$ is a nonsingular complex algebraic variety for generic $p\in\mathbb{C}^{k}$.
  • (real)
@@
    (degAllG) implies (rad), (reg) and (real).
  7. (vii)
    (flatG) implies (setZ), and the converse is true if $\mathcal{I}$ is nonsingular.
@@
For (ii), for the implication (setZ)$\Rightarrow$(dimG), Lemma 2.8 gives that $\mathcal{Z}_{\mathcal{P}}$ is Zariski dense whenever $\mathcal{Z}_{\mathcal{P}}(\mathcal{X})$ is, and hence (dimG) follows from Theorem 2.3(ii). The implication (dimG)$\Rightarrow$(dim1) is immediate. Finally, if (dim1) holds, then Theorem 2.3 gives that $\mathcal{Z}_{\mathcal{P}}$ is Zariski dense, and by Lemma 2.8 so is $\mathcal{Z}_{\mathcal{P}}(\mathcal{X})$, giving (setZ).
@@
For (vi), the implications from (degAllG) to (real) and (reg) follow from Proposition 2.11. The implication (degAllG)$\Rightarrow$(rad) follows from Proposition 2.16.
Finally, (vii) is the content of Proposition 2.5. ∎
@@
is irreducible, Theorem 2.3 holds also in the affine setting over $\mathbb{C}$ (rather than $\mathbb{C}^{*}$) after replacing $\mathbb{V}_{\mathbb{C}^{*}}(F_{p})$ by $\mathbb{V}_{\mathbb{C}}(F_{p})$. Additionally, Lemma 2.8, Propositions 2.11, 2.14 and 2.16, as well as Theorem 2.15 extend easily to $\mathbb{C}$ by allowing $\mathcal{X}\subseteq\mathbb{C}^{n}$. We conclude that Theorem 2.17 holds in the affine setting over $\mathbb{C}$ as long as $\mathcal{I}_{\mathbb{C}}$ is irreducible and $\mathcal{I}_{\mathbb{C}}\cap(\mathcal{P}\times\mathcal{X})$ is Zariski dense in $\mathcal{I}_{\mathbb{C}}$.
In Section 3, we will see that for augmented vertically parametrized systems, the nine first statements (deg1)–(dimG) from the beginning of the subsection are equivalent. Thus, all the desired properties of the polynomial system will rely on the existence of a nondegenerate zero in the complex torus. The key part of the argument will be to show that $\mathcal{Z}$ has nonempty Euclidean interior if and only if $F$ has a nondegenerate zero. This phenomenon need not happen for general families, as example (1.8) shows.
@@ 2.6. Zariski denseness and nonempty Euclidean interior
For part (i), as $S$ is constructible, $S=\bigcup_{i=1}^{t}Z_{i}\cap U_{i}$, for some irreducible Zariski closed sets $Z_{i}\subseteq\mathbb{C}^{k}$ and nonempty Zariski open subsets $U_{i}\subseteq\mathbb{C}^{k}$. The statement now follows from the fact that the Zariski closure satisfies $\overline{S}=\bigcup_{i=1}^{t}Z_{i}$. In case (ii), [BCR98, Proposition 2.8.2] gives that $S$ is Zariski dense in $\mathbb{C}^{k}$ if and only if its semialgebraic dimension is $k$, which in turn is equivalent to $S$ having nonempty Euclidean interior in $\mathbb{R}^{k}$. ∎
@@ 3.3. Nondegenerate zeros and generic consistency
Let $G\in\mathbb{C}^{m\times(m-s)}$ be Gale dual to $C$, in the sense that the columns of $G$ form a basis for $\ker(C)$. The parametrization $\phi$ in (3.3) gives rise to a parametrization of $\mathcal{Z}$ via the Laurent polynomial map
@@
As $\mathcal{D}=\Phi(\Delta)$ by Proposition 3.2(iii), it follows from Proposition 3.3 that $\mathcal{D}$ is the set of critical values of $\Phi$. Now Sard’s Lemma (see, e.g., [SEDS17, Proposition B.2] and [Mum76, Proposition 3.7]), adapted to very affine varieties, gives that $\dim(\overline{\mathcal{D}})<m+\ell$. ∎
@@
For $\mathbb{G}=\mathbb{R}^{*}$ and $\mathbb{G}=\mathbb{R}_{>0}$, it is straightforward to verify that restricting the map $\phi\colon\ker(C)\times(\mathbb{C}^{*})^{n}\to\mathcal{I}$ to $(\ker(C)\cap\mathbb{G}^{m})\times\mathbb{G}^{n}$ provides a parametrization of $\mathcal{I}\cap((\mathbb{G}^{m}\times\mathbb{R}^{\ell})\times\mathbb{G}^{n})$, which after projection gives a parametrization of $\mathcal{Z}_{\mathbb{G}^{m}\times\mathbb{R}^{\ell}}(\mathbb{G}^{n})$.
@@ 3.4. The main theorem on generic dimension
  1. (i)
    The statements (deg1), (degX1), (degXG), (degAllG), (setE), (setZ), (flatG), (dim1) and (dimG) are all equivalent to the condition
@@
  2. (ii)
    Any of the statements mentioned above implies (dimX), (reg) and (real).
  3. (iii)
@@
The equivalence between (deg1) and the matrix rank condition is Proposition 3.2(ii). As $\mathcal{I}$ is irreducible and nonsingular of dimension $m+n-s$, the assumptions for Theorem 2.17 are satisfied, and there is an equivalence between (setZ) and (flatG). Therefore, for (i) and (ii), it suffices to show that (setZ) implies (degAllG) and (setE). The equivalence between (setZ) and (setE) is Proposition 2.22(ii). By Lemma 2.8 and Proposition 2.22, $\mathcal{Z}$ has nonempty Euclidean interior in $\mathbb{C}^{m}\times\mathbb{C}^{\ell}$ if (setZ) holds. By Theorem 3.5, this implies (degAllG). Finally, to prove (iii), we note that if (degAllG) holds, so does (rad) by Theorem 2.17. If, on the other hand, (degAllG) does not hold, then (i) gives that (setZ) does not hold either, and hence it follows by Theorem 2.3 that $\mathbb{V}_{\mathbb{C}^{*}}(F_{a,b})=\varnothing$ for generic $(a,b)\in\mathbb{C}^{m}\times\mathbb{C}^{\ell}$. Thus, generically, $\langle F_{a,b}\rangle=\mathbb{C}[x^{\pm}]$, which is radical. ∎
@@
    If (degAllG) holds under the assumptions of Theorem 3.7, then we also have that all zeros of $F_{a,b}$ are nondegenerate for generic $(a,b)\in\mathcal{Z}_{\mathcal{P}}(\mathcal{X})$.
@@
    When $\ell=0$, the rank condition in Theorem 3.7(i) simplifies to
@@
Theorem 3.7 tells us that for augmented vertical systems, only one of two extreme scenarios occurs, provided the parameter sets $\mathcal{A}$ and $\mathcal{B}$ and the domain $\mathcal{X}$ of the variables satisfy some mild conditions: Either the system is generically consistent, has generically the expected dimension $n-(s+\ell)$, and generically all zeros are nondegenerate, or the system is generically inconsistent, has never the expected dimension over $\mathbb{C}$, and all zeros are degenerate.
@@
The conclusions of Theorem 3.7 also hold for parametric systems $F_{\cdot,b}$ given by the restriction of an augmented vertical system $F$ to a fixed value of $b\in L((\mathbb{C}^{*})^{n})$. In this case, the incidence variety is irreducible and nonsingular of dimension $n+m-(s+\ell)$, with parametrization
@@ 3.5. Computational considerations
The rank condition from Theorem 3.7 can be checked in the following way: Let $G\in\mathbb{C}^{m\times(m-s)}$ be a Gale dual to $C$, whose columns form a basis for $\ker(C)$. We want to check whether there exists some $(u,h)\in\mathbb{C}^{m-s}\times(\mathbb{C}^{*})^{n}$ such that
@@ 3.6. Freely parametrized systems
We now turn our attention to the special case of freely parametrized systems, and record some corollaries of our main results.
Given finite sets $\mathcal{S}_{1},\dots,\mathcal{S}_{s}\subseteq\mathbb{Z}^{n}$, we consider the corresponding freely parametrized family
@@
and say that a property holds generically in $\mathcal{F}_{\mathrm{free}}(\mathcal{S}_{1},\dots,\mathcal{S}_{s})$ if it does so under the isomorphism $\mathcal{F}_{\mathrm{free}}(\mathcal{S}_{1},\dots,\mathcal{S}_{s})\cong\prod_{i=1}^{s}\mathbb{C}^{\mathcal{S}_{i}}$ that identifies each polynomial with its coefficients.
@@
Similarly, let $M=[M_{1}\ \cdots\ M_{s}]\in\mathbb{Z}^{n\times m}$ be the block matrix where the columns of $M_{i}\in\mathbb{Z}^{n\times m_{i}}$ are the elements of $\mathcal{S}_{i}$ in some fixed order. For the vertical system $F=C(a\star x^{M})$, it then holds that $\mathcal{F}_{\mathrm{free}}(\mathcal{S}_{1},\dots,\mathcal{S}_{s})=\{F_{a}:a\in\mathbb{C}^{m}\}$. We refer to this $F$ as the vertical system associated with $\mathcal{S}_{1},\dots,\mathcal{S}_{s}$.
Let $V_{1},\ldots,V_{s}\subseteq\mathbb{R}^{n}$ be sets of vectors. Following the terminology in [Per69, Yu16], we define an independent transversal of $(V_{1},\ldots,V_{s})$ to be a linearly independent tuple $(v_{1},\ldots,v_{s})\in\prod_{i=1}^{s}V_{i}$. It is a well-known linear algebra fact (see, e.g., [Per69, Theorem 1] and [Kho16, Theorem 4]) that the existence of such an independent transversal is equivalent to
$\dim\!\Big{(}\sum_{j\in J}\operatorname{span}_{\mathbb{R}}(V_{i})\Big{)}\leq|J|\quad\text{for all $J\subseteq[s]$}.$ (3.10)
Consider now support sets $\mathcal{S}_{1},\ldots,\mathcal{S}_{s}\subseteq\mathbb{Z}^{n}$, and let
$\operatorname{Lin}(\mathcal{S}_{i}):=\operatorname{span}_{\mathbb{R}}\{u-v:u,v\in\mathcal{S}_{i}\}\subseteq\mathbb{R}^{n}$
denote the direction of the affine hull of $\mathcal{S}_{i}$. Then $(\mathcal{S}_{1},\ldots,\mathcal{S}_{s})$ is called an essential family in [Stu94, BS24] if $(\operatorname{Lin}(\mathcal{S}_{1}),\ldots,\operatorname{Lin}(\mathcal{S}_{s}))$ satisfies (3.10). It is shown in [Yu16, Lemma 1] and [Kho16, Theorem 11] that this completely characterizes when $\mathcal{F}_{\mathrm{free}}(\mathcal{S}_{1},\ldots,\mathcal{S}_{s})$ is generically consistent. We recover this fact as a corollary of Theorem 3.7.
@@
Let $\mathcal{S}_{1},\dots,\mathcal{S}_{s}\subseteq\mathbb{Z}^{n}$, and let $\mathcal{G}_{i}\subseteq\mathbb{C}^{n}$ be a generating set of $\operatorname{Lin}(\mathcal{S}_{i})$ for $i=1,\dots,s$. The following statements are equivalent:
@@
    The tuple $(\mathcal{G}_{1},\ldots,\mathcal{G}_{s})$ admits an independent transversal.
@@
    The tuple $(\operatorname{Lin}(\mathcal{S}_{1}),\ldots,\operatorname{Lin}(\mathcal{S}_{s}))$ admits an independent transversal.
@@
$C_{i}\operatorname{diag}(G_{i}u_{i})=[u_{i1}\quad\dots\quad u_{i,m_{i}-1}\quad-(u_{i1}+\dots+u_{i,m_{i}-1})].$
@@
The specialization of Theorem 3.7 for vertical systems to the case $\mathcal{A}=\mathbb{R}^{m}_{>0}$ and $\mathcal{X}=\mathbb{R}^{n}_{>0}$ allows us to study freely parametrized systems with real coefficients with prescribed sign. Specifically, we consider maximal dimensional orthants $\mathcal{O}_{i}$ of $(\mathbb{R}^{*})^{\mathcal{S}_{i}}$, and consider the subset $\mathcal{F}_{\mathcal{O}_{1},\dots,\mathcal{O}_{s}}(\mathcal{S}_{1},\dots,\mathcal{S}_{s})$ of $\mathcal{F}_{\mathrm{free}}(\mathcal{S}_{1},\dots,\mathcal{S}_{s})$ consisting of the polynomials in the image of $\prod_{i=1}^{s}\mathcal{O}_{i}$ under the isomorphism $\prod_{i=1}^{s}\mathbb{C}^{\mathcal{S}_{i}}\cong\mathcal{F}_{\mathrm{free}}(\mathcal{S}_{1},\dots,\mathcal{S}_{s})$.
@@
Let $\mathcal{S}_{1},\dots,\mathcal{S}_{s}\subseteq\mathbb{Z}^{n}$. For $i=1,\dots,s$, let $\mathcal{G}_{i}\subseteq\mathbb{C}^{n}$ be a generating set of $\operatorname{Lin}(\mathcal{S}_{i})$ and $\mathcal{O}_{i}$ be a maximal dimensional orthant of $(\mathbb{R}^{*})^{\mathcal{S}_{i}}$. The following statements are equivalent:
@@
    The semialgebraic set $\mathbb{V}_{\mathbb{R}^{*}}(F)\cap\mathbb{R}^{n}_{>0}$ has dimension $n-s$ for $F$ in a subset of
    $\mathcal{F}_{\mathcal{O}_{1},\dots,\mathcal{O}_{s}}(\mathcal{S}_{1},\dots,\mathcal{S}_{s})$ with nonempty Euclidean interior.
@@
    The tuple $(\mathcal{G}_{1},\ldots,\mathcal{G}_{s})$ admits an independent transversal and $\mathcal{O}_{i}\notin\{\mathbb{R}^{\mathcal{S}_{i}}_{>0},\mathbb{R}^{\mathcal{S}_{i}}_{<0}\}$ for all $i=1,\dots,s$.
@@
In the freely parametrized setting, it is known from [BS24, Proposition 1.12] that one can obtain the generic properties predicted by Corollary 3.14, by only requiring the coefficients corresponding to the vertices of the Newton polytopes to be generic (while the other coefficients can be arbitrarily fixed). It is an interesting problem for future research to determine which parameters can be fixed in a vertical system while still retaining the generic properties given by Theorem 3.7.